OpenStudy (anonymous):

factor (x+9)^-1/3-(x+9)^-2/3 plz helpp!

OpenStudy (anonymous):

\[\frac{x+8}{3 (x+9)^2} \]

jhonyy9 (jhonyy9):

\[(x+9)^{-1/3}-(x+9)^{-2/3}=1/(x+9)^{1/3} - 1/((x+9)^{2})1/3 =\]

jhonyy9 (jhonyy9):

\[((x+9)-1)/((x+9)^{2})^{1/3}\]

jhonyy9 (jhonyy9):

\[(x+8)((x+9)^{2})^{1/3}/(x+9)^{2}\]